Recent Log Entries

Hall A Target Change 10 uA of CW on Carbon Hole

Log Number: 4186862Logbook(s): Entry Makers: emonkBody: 

Hall A Target change request -

Target: Carbon Hole with 10 uA of CW

Figure 1: Target Ladder (Before Adjustments)
Figure 2: Ion Chambers adjusted
Figure 3: Target Ladder (After Adjustments)
Figure 4: Max Juice Settings
Figure 5: Hall A Operational Restrictions
Figure 6: Beam Authorization

Hall A No Target IC Calibration up to 10uA

Log Number: 4186861Logbook(s): Entry Makers: moserBody: 

Figure 1: Setup, raster on at nominal settings along with SBS dipole and correctors. BigBite also on at 750A.
XSOFs for no target, i.e. PolHe3 cell, dialed in.

Figure 2: Orbit. Figures 3-6: iCalibrate setup.

Figure 7: Plots during calibration.

Figures 8-14: Plots. Only 1H05, target upstream, has a better linear than log fit.

A-2023-09-25_0123_No_Target_Pass-4.hcd in ~mccops/IonChambers/HallA_2023/

iCalibrate: Hall A

Log Number: 4186858Logbook(s): Body: Calibration ParametersHall:ATarget: Carbon OpticsPass: 4Note: Solid Thin for PolHe3Calibrated On:2023-Sep-14 04:42Calibrated By: moserFile:A-2023-09-14_0442_Solid_Thin_Targets_like_Carbon_Optics_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

Checking the Phase for 1L26-3

Log Number: 4186857Logbook(s): Entry Makers: emonkBody: 

Using the 1A SLM to check the phase of 1L26-3. Sending about 5 uA of tune beam with the A laser to manually adjust the PSET value, and noting the beam displacement (checking for optimal/peak energy).

The cavity didn’t require any adjustments in phase value.

Initial PSET value: -93.90
Final PSET value: -93.90

Initial PMES value: 39.5013
Final PMES value: 39.5068

Both PSET and PMES values measured in degrees.

Autolog save

Log Number: 4186850Logbook(s): Tags: Body: saver ran from opsl03

Saved: emonk 09-25-23 00:57:36 {A:4 B:5 C:5 E:1047 A=CW B=BS C=CW D=BS 0R08=39.75 No dumps in.

HYST AREA 6-3-2 Cycles Magnets

Log Number: 4186846Logbook(s): Body: 

HYST AREA Version 6-3-2 has successfully cycled the following:

MBD1C00V
MBD1C00AV
MBC1C01H
MBC1C02V
MBC1C03H
MBC1C04H
MBC1C05V
MBC1C06H
MBC1C07H
MBC1C07V
MBC1C09V
MBC1C11V
MBC1C13V
MBC1C15V
MBC1C18H
MBC1C18V
MBC1C20H
MBC1C20V
MBT1P01H
MBT1P04H
MBC1H01H
MBC1H01V
MBD1H04V
MBD1H04H

Total Time: 2.433 mins

HYST AREA 6-3-2 Cycles Magnets

Log Number: 4186845Logbook(s): Body: 

HYST AREA Version 6-3-2 has successfully cycled the following:

MBD1C00V
MBD1C00AV
MBC1C01H
MBC1C02V
MBC1C03H
MBC1C04H
MBC1C05V
MBC1C06H
MBC1C07H
MBC1C07V
MBC1C09V
MBC1C11V
MBC1C13V
MBC1C15V
MBC1C18H
MBC1C18V
MBC1C20H
MBC1C20V
MBT1P01H
MBT1P04H
MBC1H01H
MBC1H01V
MBD1H04V
MBD1H04H

Total Time: 2.383 mins

Addressing Hall A 1H05 and 1H04B Ion Chamber Trips -- 8E02.X position adjustment silver bullet

Log Number: 4186843Logbook(s): Entry Makers: moserBody: 

When trying to deliver CW on Carbon Hole in Hall A, the first CW after a 2+ day down, found we couldn't deliver any appreciable beam current without 1H05, target upstream, and 1H04B, collimator, ion chamber trips. Switching to tune, could even trip these in tune-mode.

Adjusted some end of the line steering -- 1C20->1H04E -- around some to see if maybe we had a bad incoming angle that was causing us to scrape the collimator, tripping ICs. Was able to get some reduction in IC signals in tune-mode, but not enough for CW delivery without IC trips.

Injector Transmission with Only 30uA to xC Hovering Around ~70%

Log Number: 4186833Logbook(s): Entry Makers: moserBody: 

Haven't directly associated the poor transmission with any issues in the machine as of yet. Perhaps it explains NPS ion chamber in Hall C spikes and general signal noise, Figure 4.

For now, won't address, just noting.

Operator SIGN IN

Log Number: 4186831Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Howard
DATE:             09/25/2023
SHIFT:               Owl

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Operator SIGN IN

Log Number: 4186829Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Blackshaw
DATE:             09/24/2023
SHIFT:               Owl

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

iCalibrate: Hall A

Log Number: 4186809Logbook(s): Body: Calibration ParametersHall:ATarget: Carbon OpticsPass: 4Note: Solid Thin for PolHe3Calibrated On:2023-Sep-14 04:42Calibrated By: moserFile:A-2023-09-14_0442_Solid_Thin_Targets_like_Carbon_Optics_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

iCalibrate: Hall A

Log Number: 4186806Logbook(s): Body: Calibration ParametersHall:ATarget: Carbon OpticsPass: 4Note: Solid Thin for PolHe3Calibrated On:2023-Sep-14 04:42Calibrated By: moserFile:A-2023-09-14_0442_Solid_Thin_Targets_like_Carbon_Optics_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

CREW CHIEF SIGN IN

Log Number: 4186801Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 09-24-23
Time : 23:12:03
Day : Sunday
Shift : Owl

Crew Chief : Moser
Program Deputy: Benesch
Team : A
Arm : Howard
Designated SSO: Blackshaw
LERF Operator:
Operators : Blackshaw, Howard

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

Pages